OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how multiplying and dividing factors with rational exponents is similar to multiplying and dividing factors with integer exponents. (2 points) Part 2: Give one example of multiplying factors with rational exponents. (1 point) Part 3: Give example of multiplying factors with integer exponents. (1 point)

OpenStudy (anonymous):

Part 1: The rules for multiplying factors with integer exponents is x^m * x^n = x^(m+n) The rules for dividing factors with integer exponents is x^m/x^n = x^(m-n). The rules are the same if the exponents are rational numbers such as a/b or c/d Example x^1/2 * x ^1/3 = x^5/6. The only difference is that you must insure erac exponent has a common denominator.

OpenStudy (anonymous):

Can you do part 2 and 3 by yourself now? @LaVidaEsGran

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

No problem!